Advanced Accounting Ch 1 Q 5

On January 1, Year 9, Harold, Inc., purchased 75,000 shares of Sally Company common stock for $1,480,000, giving Harold 25 percent ownership and the ability to apply significant influence over Sally.  Any excess of cost over book value acquired was attributed solely to goodwill.

    Sally reports net income and dividends as follows.  These amounts are assumed to have occurred evenly throughout these years.  Dividends are declared and paid in the same period.

 

                     

                                                                                                                   Annual

                                                                                                            Cash Dividends

                                                         Net Income                               (paid quarterly)

                         Year 9                        $340,000                                        $120,000

                         Year 10                        480,000                                          140,000

                         Year 11                        600,000                                          160,000

 

   On July 1, Year 11, Harold sells 12,000 shares of this investment for $25 per share, thus reducing its interest from 25 to 21 percent, but maintaining its significant influence.

   Determine the amounts that would appear on Harold’s Year 11 income statement relating to its ownership and partial sale of its investment in Sally’s common stock.
Answer #1

Step 1: Find net income for each year.

Year 9:  340,000 x 25% = 85,000

Year 10: 480,000 x 25% = 120,000

Year 11 (first half) : 600,000 x 25% x 50% = 75,000

Year 11 (second half): 600,000 x 21% x 50% = 63,000


Step 2: Find dividends for each year

Year 9: 120,000 x 25% = 30,000

Year 10: 140,000 x 25% = 35,000

Year 11 (first half): 160,000 x 25% x 50%

Year 11 (second half): 160,000 x 21% x 50% = 16,800


Step 3: Find the balance of net income and dividends.(Add the net incomes and subtract the dividends.) 

1,480,000 common stock purchased + 85,000 + 120,000 + 75,000 - 30,000 - 35,000 - 20,000 = 1,675,000 book value


Step 4: Compare the fair value to the book value.

Fair value = 12,000 cash x 25 per share = 300,000

Book value = 1,675,000 / 75,000 purchased x 12,000 shares sold = 268,000 

Fair value - book value = gain 

300,000 - 268,000 = 32,000 gain on sale


For the equity in earnings of Affiliate, we would add both halves of Year 11's net income. 138,000 =  75,000 first half + 63,000 second half 


Income Statement Effects

Equity in Earnings Affiliate = 138,000

Gain on Sale = 32,000










answered by: anonymous